OpenStudy (emmynimmy):

Solve the quadratic equation by factoring. (Enter your answers as a comma-separated list.) x2 + 2x − 35 = 0 I keep getting -7 or 5 and it is incorrect, what is the correct answer? I have tried everything.

OpenStudy (jtm-pa):

(x+7)(x-5)

OpenStudy (jtm-pa):

though when you take the zeros of that equation it gives you the solutions -7, 5

OpenStudy (emmynimmy):

it says enter as comma separated, what do i do?

OpenStudy (emmynimmy):

@jtm-PA

OpenStudy (jtm-pa):

:\ im not sure exactly I would say what you wrote was correct considering those are the zeros of the function.. maybe you should contact your teacher

OpenStudy (emmynimmy):

Solve the quadratic equation by factoring. (Enter your answers as a comma-separated list.) 4x^2 = 9

OpenStudy (emmynimmy):

@jtm-PA

OpenStudy (jtm-pa):

\[x \pm \frac{ 3 }{ 2 }\]

OpenStudy (emmynimmy):

is there just one? or two?

OpenStudy (jtm-pa):

2

OpenStudy (emmynimmy):

what is the second one? is there a calculator online that i can search these with?

OpenStudy (jtm-pa):

um it is postive 3/2 and negative 3/2

OpenStudy (emmynimmy):

I have a lot more of these. is there an online calculator to solve these?

OpenStudy (emmynimmy):

The quadratic, (2x2 − 20x + 48), factors as (2x − 8)(x − 6). (a) Is the expression completely factored? No (b) If not, factor completely. (If it is already in fully factored form, restate the factored form.)

OpenStudy (jtm-pa):

I guess wolframalpha.com is a pretty good calculator

OpenStudy (jtm-pa):

for that equation the answer is yes it is fully factored

OpenStudy (emmynimmy):

no the answer is no i already checked

OpenStudy (twerk):

I think it's x= 35/4 or Decimal I got x=8.75